OpenStudy (anonymous):

Mr. Steward has a jar in which each of his 26 students has placed a stick w/his or her name on it. If there are 10 girls in the class, what is the probability that a boys name will be drawn if the teacher only pulls one stick?

OpenStudy (anonymous):

\[P=\frac{26-10}{26}=0.615\](rounded)

OpenStudy (anonymous):

well assuming that if they arent girls then they are boys first we find the amount of boys there are 26-10=16 so there are 16 boys then we place the number of desired out comes(16) over number of possible out comes(26) P(B)=16/26 reduce P(B)=8/13

OpenStudy (anonymous):

wait why would u reduce???

OpenStudy (anonymous):

to get the fraction in lowest form. or u could put it as a decimal like eseidl did